Your broker has recommended that you purchase stock in National Bank & Trust, Inc. National Bank & Trust recently paid its annual dividend of $8. The firm has an ROE of 15% and pays out 50% of its earnings as dividends. Based on your analysis, you estimate that the stock has a required return of 15%. What is the intrinsic value of this stock?

Group of answer choices

$119.60

$123.32

$114.67

$106.98

$112.58
Answer #1

DoE 8 g= RoEx Retention ratio こ • 15 X (1-050) 7.51. Re = 15% In hinsic value = Doll+g) Re- g 8 x (1 + 7.500) •15 - 075 = $ 1